OpenStudy (anonymous):

Having some conceptual trouble with this problem: "A falling object travels one-fourth of its total distance in the last second of its fall. What height was it dropped from?" Would someone please help me set this one up?

OpenStudy (jamesj):

So suppose the object is dropped from rest at a height of h. You know that h = (1/2)gT^2, from basic kinematics (yes?). T is the time for the object to reach the zero mark and T is the time taken; g the acceleration due to gravity. Now, the new information you are given in the problem is that the distance travelled in the last second is h/4. What is that distance travelled in the last second? Well, at time t = T-1, the object has falled a distance of d1 = (1/2)g(T-1)^2. At time t = T, we know the distance travelled is d2 = (1/2)gT^2 = h. Hence in the last second, the object has travelled a distance of d2 - d1 = h - (1/2)g(T-1)^2 and you know that is equal to h/4 You will now have to manipulate these equations to eliminate T from them in order to express h purely in terms of g.

OpenStudy (jamesj):

...or solve for T in terms of g. Either way!

OpenStudy (anonymous):

I'm with you so far... do I eliminate T by using some other form of T?

OpenStudy (jamesj):

Well, play around with it and see what you can do. I think I would take this last expression h - (1/2)g(T-1)^2 and set it equal to h/4 and see what that tells me about T. Remember that we can also write h = (1/2)gT^2

OpenStudy (anonymous):

Oh, I see! By the time h/4 = h - (1/2)g(T-1)^2 is solved for T, I'll have a new expression to plug in for T.

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

Thanks so much. That cleared it right up :)

OpenStudy (jamesj):

What answer do you get for h in terms of g?

OpenStudy (anonymous):

Actually... I've tried it a few different ways now and something still seems wrong. What you're saying makes sense, but I'm getting just strange answers, like t = sqrt(h/10g)... where did I go wrong?

OpenStudy (anonymous):

Or \[{3h \over g }=t^2+2t-1\]

OpenStudy (jamesj):

Yes, I think you get a quadratic expression. The answer isn't so neat and elegant.

OpenStudy (jamesj):

Be careful with signs btw.

OpenStudy (anonymous):

Orginally I got \[-{3h \over g} = -t^2-2t+1\] Should I leave it like that?